The mesa Redbirds football team plays in a stadium with seating capacity of 80,000. However, during the past season, attendance averaged only 50,000. The average ticket price was $30. If the price elasticity is -4, what would the team have to charge in order to fill the stadium?If the price were to be decreased to $27 and the average attendance increased to 60,000, what is the price elasticity?
1
Expert's answer
2018-11-28T11:03:10-0500

The point price elasticity of demand is calculated as


K_d=∆Q/Q1:∆Р/Р1

So,

∆P=(∆Q×P1)/(Q1×K_d )


∆P=((80,000-50,000)×$30)/(50,000×(-4) )=4.5


So, the team has to charge the price 30-4.5=$25.5


If the price were to be decreased to $27 and the average attendance increased to 60,000, the price elasticity is


K_d=∆Q/Q1:∆Р/Р1=(60,000-50,000)/50,000÷(27-30)/30=-2